OpenStudy (anonymous):

what is the formula used to find the square inch of a 9 inch pizza?

OpenStudy (anonymous):

I'm guessing you want the area of the pizza? Basically a 9-inch pizza is the diameter of the pizza. So you know that R (Radius) = D/2 = 9/2 In this case = 4.5 Inches Then area of a circle (A pizza is a circle!) is given by: \[A = \pi r^2\] So just plug in R for that equation, and then you get area of the circle, which is in square inches.